i am trying to find out how you would find the super compaonent of a full package of $85000.
I thought you would do:
xy=z
x=unknown
y=.09
z=85000
so what is x? and how do you do it and an algebra formula.

Answer by venugopalramana(3286)

XY=Z....................I
Z=85000....................II
Y=0.09.............................III
SUBSTITUTING II AND III IN I
X*0.09=85000
X=85000/9=9444.44